Solve the equation.
6/m-5 - 2/m+5 = 8/m^2-25

Answer by lwsshak3(11628)

Solve the equation.
6/m-5 - 2/m+5 = 8/m^2-25
m^2-25=(m+5)(m-5)
LCD:(m+5)(m-5)
6(m+5)-2(m-5)=8
6m+30-2m+10=8
4m=-32
m=-8